OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 2. Find x when y = 12.

OpenStudy (blurbendy):

y = k/3x 2 = k/3(6) 2 = k / 18 k = 36 12 = 36/x x = 3

OpenStudy (anonymous):

thanks!